When the temperature of a gas is raised from 27`c to 90`c the percentage increase in the ram velocity of molecular

Grade:11

1 Answers

Arun
6 years ago
Vrms = √(3RT / Mo)
i.e. Vrms ∝ √T
T1 = 27°C = 27 + 273 = 300K
T2 = 90°C = 90 + 273 = 363K
∴  {Vrms2 / Vrms1} = √(T2 / T1) = √(363 / 300) = 1.1
% increase in rms velocity = {(Vrms2 – Vrms1) / (Vrms1)} × 100
= {(1.1 – 1) / 1} × 100
= 0.1 × 100
% increase in rms velocity = 10%